A goldsmith has 2 alloys that are different purities of gold.
the first is three-fourths pure gold and the second is five-twelfths pure gold. how many ounces of each should be melted and mixed in order to obtain a 6 oz. mixture that is two-thirds pure gold?

Answer by stanbon(75887)

A goldsmith has 2 alloys that are different purities of gold.
the first is three-fourths pure gold and the second is five-twelfths pure gold. how many ounces of each should be melted and mixed in order to obtain a 6 oz. mixture that is two-thirds pure gold?
----------------------------
Equation:
gold + golf = gold
(3/4)x + (5/12)(6-x) = (2/3)6
---
Multiply thru by 12 to get:
9x + 5*6 - 5x = 8*6
4x = 3*6
x = 9/2 = 4.5 oz (amt of 3/4th pure gold needed for the mixture)
--------------------
6-4.5 = 1.5 oz (amt of 5/12th pure gold needed for the mixture
=============================
Cheers,
Stan H.